OpenStudy (anonymous):

Wynne and Emmett are brother and sister. Wynne runs 180 yards Emmett 120 yards how old are they? how far they run is based on age ratio to each other.

OpenStudy (briannaontimeforreal):

i can help you. so if you are tryna find the age of some one you can multiply so 180 times 120 equals 21600 and there is your answer

OpenStudy (anonymous):

not that big they cant be that old

OpenStudy (anonymous):

they are kids

OpenStudy (ghostgate):

Their ratio bases off of 3:2. 3 representing the brother, 2 representing the sister. I guess the way it works is that the brother is 1.5 times the sister's age, we can conclude that the sister's age can be represented by 'x' whereas the brother is represented as '1.5', together they ran a total of 300 yards, so we can use this to set up the problem: x + 1.5x = 300 - Now we add like terms: 2.5x = 300 - Divide. x = 120 - Of course upon adding that back in we just reroute back to only knowing how far they ran. Hmmm... The problem is this, we have no idea what the variable is to determine their ages, the ratio is 3:2, but... the only solutions I can think of are: If 10 yards = 1 Year of age(roughly) we have: 18 = Brother 12 = Sister If 20 yards = 1 Year of age(roughly) we have: 9 and 6. Regardless I can't say for certain what their ages are. :/ I hope this helps in someway! Have a great day! ;) {---Ghostgate---}